please help me do this!!!!
(a) graph x^2/16 - y^2/25 = 1 show how you arrived at the graph by determining the (b) x-intercepts, (c) extent of the graph and the (d)asymptotes.
(a)_____________
(b)_____________
(c)_____________
(d)_____________

Answer by stanbon(48569)

a) graph x^2/16 - y^2/25 = 1 show how you arrived at the graph by determining the (b) x-intercepts, (c) extent of the graph and the (d)asymptotes
-------------------------
x-intercepts:
Let y = 0, then x^2 = 16; x=4 or x=-4
-------------------------
y-intercept:
Let x = 0, then y^2 = -1; y is imaginary so no y intercepts
--------------------
Extent of the graph:
It is a hyperbola with center at the origin; vertices are the x-intercepts;
opening to the left and to the right.
-------------------------------
Asymptotes:
a = 4 ; b = 5
y = (b/a)x = (5/4)x
y = (-b/a)x = (-5/4)x
====================
Cheers,
Stan H.